# PEVLM: Parallel Encoding for Vision-Language Models

### v1

@@ Abstract
Vision-Language Models (VLMs) have demonstrated strong performance in video-language tasks, yet their application to long video understanding remains constrained by the quadratic complexity of standard attention mechanisms. In this paper, we propose PEVLM, a parallel encoding strategy specifically designed to improve the prefill efficiency of VLMs without requiring model finetuning. PEVLM partitions the input into block-wise segments with a shared sink, preserves Full-Attention positional embeddings, and aligns attention weights to mimic Full-Attention distributions. This design reduces attention computation from $O((T\times N)^{2})$ to $O(T\times N)$ while maintaining high accuracy. Extensive experiments on the LongVideoBench benchmark show that PEVLM achieves up to 8.37% accuracy improvement over existing inference-efficient methods and delivers up to 7.47x speedup in attention computation and 40% reduction in end-to-end latency. Under strict latency constraints, PEVLM significantly outperforms baselines, raising accuracy from 23.26% to 61.03%. These results highlight PEVLM’s effectiveness for low-latency, long-context video understanding, making it well-suited for real-world applications such as autonomous driving.
1 Background
1.1 Challenges in Long Context Processing
In the realm of computer vision and natural language processing integration, Vision-Language Models (VLMs) have demonstrated remarkable performance in various multimodal tasks, particularly in video-text understanding and generation (alayrac2022flamingo, ; radford2021learning, ; vlm_performance_example, ). However, their application to long video processing remains computationally challenging due to the quadratic complexity of self-attention mechanisms (long_video_processing_challenges_example, ; vaswani2017attention, ). Specifically, VLMs usually extract frame-wise features and integrate textual context information, leading to a drastic increase in context length as each frame is tokenized into multiple embeddings.
The high computational complexity of VLMs presents a significant barrier to their practical use and deployment in real-world scenarios (li2022video, ). Take autonomous driving as an example: training advanced autonomous driving models requires extensive long video data, which in turn demands efficient data pre-processing and annotation techniques. Traditional manual annotation methods and classical annotation models are inefficient in handling the huge volume of long video data. Although VLMs offer the possibility of automated annotation, the computational cost of processing long videos is prohibitively high. Moreover, autonomous driving systems impose strict requirements on low latency execution and high accuracy due to the critical need for real-time decision making to ensure safety (zhou2022efficient, ). However, on-board computing resources are severely limited, and the latency introduced by long-context attention mechanisms in VLMs significantly degrades the response time of autonomous systems. This situation highlights the urgent challenge of improving the inference performance of the VLMs with limited latency while maintaining its accuracy.
In autonomous driving scenarios, the system typically does not generate lengthy CoT (chain-of-thought) analyses, but instead needs to output only short environmental features and final decisions. This results in the prefill stage accounting for a substantial proportion of the overall inference time, becoming the primary performance bottleneck. However, most existing methods for optimizing the prefill stage involve a trade-off between performance improvement and accuracy degradation. Given that autonomous driving and other safety-critical applications are highly intolerant of accuracy loss, these methods are not applicable. Therefore, our research focuses on search for novel strategies to optimize the prefill stage latency of VLMs with minimal impact on accuracy.
1.2 Existing Solutions and Limitations
@@
A widely adopted technique in LLMs is the parallel-encoding mechanism (li2024focusllmpreciseunderstandinglong, ; liu2024star, ; ma2025blockattentionefficientprefilling, ; ratner2023parallelcontextwindowslarge, ; yen2024longcontextlanguagemodelingparallel, ; yang2025ape, ). In this approach, the input context is divided into blocks, and each block is encoded into Key-Value (KV) states independently. This means that tokens from different blocks do not attend to each other during encodin. During decoding, however, the model computes attention using the complete KVCache that contains all historical blocks.
In addition, several block-based methods have been proposed that introduce training procedures to further enhance model accuracy and performance (ma2025blockattentionefficientprefilling, ; yuan2025nativesparseattentionhardwarealigned, ). However, not all application scenarios have sufficient resources to support large-scale training. On the other hand, many platforms offering inference services cannot control users’ preferences for model usage, nor can they mandate further fine-tuning on existing models. This is especially true for autonomous driving models that have undergone continuous training and iteration over months or even years. Therefore, we prioritize finetune-free approaches for optimization.
The parallel-encoding mechanism omits the attention calculation between blocks, reduces the computational complexity from $O(L^{2})$ to $O(L)$. It also mitigates the problem of "lost in the middle" (Liu2023LostIT, ; yang2025ape, ). Several architectural improvements have been proposed to enhance the parallel-encoding mechanism:
  • Adaptive Parallel Encoding (APE) improves prefill efficiency by introducing shared prefixes, an attention temperature parameter, and a scaling factor to better align the distribution of parallel-encoding with that of sequential encoding (yang2025ape, ). Although these adjustments lead to improved accuracy, they are highly sensitive to length, quantity, and content, making them difficult to tune and deploy in practice.
  • Star Attention, introduces a two-phase block-sparse approximation that improves computational efficiency by sharding attention across multiple hosts while minimizing communication overhead and reducing computational complexity (liu2024star, ). Although primarily designed for distributed settings, its core algorithms are mathematically similar to APE when temperature and scaling are not applied.
  • Block-Attention applies position re-encoding to allow reuse of prefilled contexts, and maintains model accuracy through fine-tuning (guu2020retrieval, ). Without fine-tuning, its accuracy suffers due to the absence of a Sink Block, though the idea of retaining Full-Attention position encodings proves insightful.
Thus, a natural research question thus arises: Can such methods be directly applied to Vision-Language Models (VLMs) to accelerate inference and deployment pipelines?
2 Our Observations and Contributions
In Table 1, we evaluate several existing methods and find that parallel-encoding-based approaches are also effective in VLMs. These methods achieve competitive accuracy on LongVideoBench (longvideobench2024, ) without requiring any modifications. This observation is consistent with findings from APE in LLM scenarios (yang2025ape, ).
2.1 Observations Specific to VLM
Despite the similarities between LLMs and VLMs, there are important distinctions. As shown in Figure 1: (i) In VLMs, attention sinks are present not only in the initial system prompt, but also in the early video frames, which aligns with the findings of Kang et al. (kang2025toldvisualattentionsink, ). To achieve optimal performance, it may be necessary to include all frames with high attention scores in the sink. This requirement may vary slightly across models; for the model shown in the figure (Qwen2.5-VL), the optimal range appears to be around 8–12 frames. (ii) Existing parallel-encoding-based are primarily designed for RAG (Retrieval-Augmented Generation) tasks. In such tasks, reusing the KVCache from historical passages is crucial for efficiency, and the order of passages often does not matter. Consequently, position embeddings are typically reused across blocks.
However, video-understanding tasks differ significantly: Maintaining the correct temporal order between video frames is essential. Reusing position embeddings indiscriminately across blocks can impair temporal comprehension of the model.
@@ 2.1 Observations Specific to VLM
(b) Reuse Position
(c) Keep Full-Attention Position Embedding
@@
Figure 1: Comparison of Attention Weight Distributions across Video Contexts. The top row shows the attention weights for all tokens, while the bottom row shows those for the video tokens only. (a) Full-Attention: The textual tokens and the initial video frames receive the highest attention weights, whereas the remaining video frames are assigned relatively lower attention weights. (b) Parallel Encoding with Reused Position Embeddings: The overall attention weights distribution resembles Full-Attention, with higher attention weights for the textual tokens and initial video frames. While compared to Full-Attention, the initial video frames receive even higher attention weights, and the remaining video frames exhibit more uniform and lower attention weights. (c) Parallel Encoding with Full-Attention Position Embedding: The attention pattern remains similar to (a) and (b), highlighting the textual tokens and the initial video frames. However, the attention weight distribution within the video frames more closely aligns with that of Full-Attention.
Therefore, as illustrated in Figure 1, we perform an attention weight analysis for VLMs and make two key findings: (i) Initial segments of both text and video receive high attention weights, emphasizing the need to include early video frames when selecting sink tokens. (ii) Using Full-Attention position embeddings (rather than reusing position embeddings across blocks) yields attention distributions more similar to Full-Attention, leading to better alignment with the model’s intended behavior. These insights guide the design of VLM-specific adaptations of parallel-encoding that preserve both efficiency and contextual integrity.
2.2 Our Contributions
Basing on these observations, we have conducted a series of studies and experiments. And proposed a novel attention mechanism for VLMs named Parallel Encoding for Vision-Language Models (PEVLM), which is a finetune-free method specifically designed to speed up prefill of VLM-based long-video processing. By alleviating computational bottlenecks in long video prefill, PEVLM facilitates the efficient deployment of VLMs in various practical application scenarios on both clouds and terminal equipments.
Our contributions involve:
  • We introduce PEVLM, a blockwise parallelization strategy for VLMs. PEVLM reduces the computational complexity of the attention mechanism from $O((T\times N)^{2})$ to $O(T\times N)$, which leading to a significant improvement in prefill efficiency.
  • We recover the accuracy of PEVLM with three alignment steps: (i) Segmenting contexts into blocks by video frames rather than tokens. (ii) Using the system promots and the initial frames of the video as Sink Block for all blocks to avoid the duplication of abnormal distribution of initial tokens. (iii) Preserving the position embeddings of Full-Attention instead of resuing position embeddings across blocks. This suggests that PEVLM serves as a performance optimization mechanism rather than endowing the model with the capability to process context lengths that exceed its inherent capacity.
  • We empirically show that (i) PEVLM gets better accuracy than Block-Attention, APE and Star Attention in long video processing tasks, achieving accuracy improvements ranging from 0.68% to 8.37% across different models. (ii) For 100k token contexts, the attention layer achieves a 7.47× speedup, while the end-to-end processing time is reduced by 40% without compromising generation quality. This highlights the practical viability and superiority of PEVLM in cloud deployment scenarios. (iii) Under fixed latency constraints, PEVLM increases accuracy from 23.26% to 61.03%, showcasing its critical value for latency sensitive applications (e.g., autonomous driving).
3 PEVLM
PEVLM is an advanced attention mechanism designed to improve computational efficiency of prefill while maintaining global contextual consistency. This approach draws from multiple existing techniques to address computational bottlenecks commonly encountered in long video processing.
3.1 Standard Attention Mechanism
@@
Attention Sink. First discovered by StreamingLLM (mit2023streamingllm, ), that one or more tokens at the start position of an LLM receive higher attention scores. Some VLM-related studies also show that the attention sink phenomenon exists in VLMs but is more distributed among image tokens (kang2025toldvisualattentionsink, ).
Position Embedding. To effectively process sequential input, LLMs require position embeddings, such as absolute position embeddings (devlin-etal-2019-bert, )(vaswani2017attention, ) and relative position embeddings (press2022trainshorttestlong, )(SU2024127063, ). For VLMs, contextual positional information is more critical. In particular, for some VLM models like Qwen2.5-VL, their position embeddings also incorporate spatio-temporal information between video frames, making position embeddings increasingly essential for VLMs.
Computational Complexity. In standard Softmax attention mechanisms, the total OPs is
$OPs_{Attn}=2HL^{2},$ (2)
where $L$ denotes the context length and $H$ is the hidden size. So, the computational complexity is $O(L^{2})$. As $L$ increases, the computational cost increases significantly, restricting the inference speed and scalability. It is even worse for VLMs. Given a video with $T$ frames, where each frame is represented by $N$ tokens (typically, $N$ ranges between 200 and 600), the resulting context length $L$ scales to $(T\times N)$ and complexity to $O((T\times N)^{2})$, making standard attention mechanisms impractical for long video understanding tasks.
3.2 Parallel Encoding
3.2.1 Partitioning Strategy
@@
APE (yang2025ape, ) and Star-Attention (liu2024star, ) have demonstrated that sharing a common attention sink across blocks significantly enhances model accuracy. Building on recent advances in efficient attention mechanisms and long-context modeling (mit2023streamingllm, ; liu2024star, ; xiao2023efficient, ; yang2025ape, ), we propose a Sink Block mechanism to improve inference quality. As illustrated in Figure 2, our partitioning strategy consists of three core components:
@@
In addition, as analyzed in Section 2.1, the number of frames selected for the Sink Block depends primarily on the distribution of attention weights. We conduct further empirical studies and analyses on this in Section 5.2. On the other hand, the number of frames per Context Block involves a trade-off between accuracy and computational efficiency, which we investigate and analyze in Section 5.3.
3.2.2 Formulations
@@
The total OPs of PEVLM is
$OPs_{PEVLMAttn}=OPs_{SinkAttn}+N\times OPs_{BlockAttn}+OPs_{QuestAttn},$ (6)
$OPs_{SinkAttn}=2HS^{2},$ (7)
$OPs_{BlocksAttn}=2HB(S+B),$ (8)
$OPs_{QuestAttn}=2HQL,$ (9)
@@
$OPs_{PEVLMAttn}=2H(S^{2}+Q^{2}+NB^{2}+QS+NQB+NSB).$ (10)
As B is a static number, the computational complexity of PEVLM is $O(L)$.
3.2.3 Position Encoding in VLMs
@@
As described in Section 2.1, similar to the case in LLMs, applying parallel-encoding-based methods in VLMs suffers from the misalignment of attention weight distributions when compared to Full-Attention. However, introducing additional hyperparameters to adjust attention weights for better alignment is not ideal for efficient deployment. Therefore, inspired by Block Attention (ma2025blockattentionefficientprefilling, ), we retain Full-Attention Position Embedding as illustrated in Figure 3.
Due to maintaining the original position embedding, PEVLM cannot enable the model to handle longer video contexts beyond its inherent capabilities like some other parallel-encoding-based methods. For example, LongVILA-7B-256f can originally process 256 frames at most. Although using PEVLM can accelerate the inference speed, it still processes 256 frames at most. This is not a problem as our goal is to accelerate the prefill.
4 Experiments
In this section, we test PEVLM from multiple perspectives against several different VLMs. The primary objective is to evaluate the accuracy, computational efficiency, and real-world applicability of PEVLM for long-video processing.
4.1 Accuracy Evaluation
The main objective of this experiment is to analyze the impact of PEVLM on accuracy in long-video understanding tasks.
4.1.1 Setup
We have chosen the LongVideoBench (longvideobench2024, ) as a benchmark, which is designed to comprehensively evaluate large multimodal models on long-term multimodal understanding. We have carried out experiments using LLaVA-Video (zhang2024videoinstructiontuningsynthetic, ), Qwen2.5-VL (qwen2.5-VL, ), and LongVILA (chen2024longvilascalinglongcontextvisual, ) as test models. To minimize any adverse effects on model accuracy, all evaluations were performed using bf16 precision.
As shown in Table 1, we compared the accuracy of Full-Attention, Block-Attention, APE, Star-Attention, and PEVLM in the LongVideoBench dataset. Since the other methods are finetune-free, we did not finetune the weights in the Block Attention test. For APE and Star-Attention, we tested two scenarios: (i) reusing position encoding across all blocks; (ii) retaining the position encoding of Full-Attention. For PEVLM, we set the sink size and block size to about 16 frames, which corresponds to approximately 4k tokens for Qwen2.5-VL and LongVILA, and about 3k tokens for LLaVA-Video.
4.1.2 Results
Table 1: LongVideoBench Accuracy
model method sink(anchor)_size block_size Correct $\uparrow$
Full Attn - - 60.43%
Block Attn (wo finetune) 0 4096 54.23%
APE 256 4096 NaN
Qwen2.5-VL- APE (retain pos) 256 4096 60.21%
7B-Instruct Star Attn 4096 4096 NaN
Star Attn (retain pos) 4096 4096 61.41%
PEVLM sys+16frames 16frames 62.23%
Full Attn - - 52.21%
Block Attn (wo finetune) 0 4096 51.76%
APE 256 4096 51.68%
LongVILA- APE (retain pos) 256 4096 53.18%
7B-256f Star Attn 4096 4096 46.75%
Star Attn (retain pos) 4096 4096 53.63%
PEVLM sys+16frames 16frames 53.70%
Full Attn - - 59.99%
Block Attn (wo finetune) 0 4096 55.50%
APE 256 4096 58.41%
LLaVA-Video- APE (retain pos) 256 4096 58.49%
7B-Qwen2 (128) Star Attn 4096 4096 58.71%
Star Attn (retain pos) 4096 4096 58.49%
PEVLM sys+16frames 16frames 59.01%
The experiments demonstrate that PEVLM achieves an accuracy comparable to or even exceeding that of Full-Attention across the three tested models. Both APE and Star-Attention can also approach or surpass Full-Attention’s accuracy when retaining Full-Attention’s position encoding. However, for Block-Attention, due to the lack of resources for alignment fine-tuning, its accuracy shows a significant decline compared to Full-Attention.
Without keeping the position encoding of Full-Attention, the output of the Qwen2.5-VL-7B-Instruct model exhibited severe output garbling and semantic confusion, and the LongVILA model also experienced noticeable performance degradation. However, the LLaVA-Video model showed insensitivity to changes in position encoding, with negligible accuracy variation observed regardless of whether Full-Attention position encoding was retained.
4.2 Performance Evaluation

### v2

@@ Abstract
Vision-Language Models (VLMs) have demonstrated strong capabilities in multimodal understanding and generation tasks. However, their application to long video understanding remains hindered by the quadratic complexity of standard attention mechanisms. In this work, we introduce PEVLM, a fine-tuning-free parallel encoding method designed to enhance the prefilling efficiency of VLMs in long video scenarios. PEVLM partitions the input video into context blocks with a shared sink block, while preserving sequential position embeddings to align the attention weight distribution with that of Full-Attention. This design reduces attention complexity from $O((T\times N)^{2})$ to $O(T\times N)$ where $T$ is the number of frames and $N$ the number of tokens per frame, without sacrificing accuracy. Extensive experiments across multiple state-of-the-art models and benchmarks demonstrate that PEVLM consistently outperforms existing parallel encoding approaches, achieving up to 7.47x speedup in attention computation and reducing end-to-end latency by 40%. Remarkably, PEVLM not only maintains high accuracy, but in some settings even surpasses Full-Attention performance. Under strict latency constraints, it achieves substantial gains, improving accuracy from 23.26% to 61.03%. These results underscore the effectiveness of PEVLM for low-latency, long-context video understanding, making it a promising solution for real-world applications.
Introduction
In recent years, Vision-Language Models (VLMs) have emerged as a central research focus at the intersection of computer vision and natural language processing, demonstrating impressive performance in multimodal understanding and generation tasks (Alayrac et al. 2022; Li et al. 2020; Radford et al. 2021; Zhu et al. 2023). As the capabilities of these models continue to strengthen, their application scenarios have also expanded, including robotics (Black, Lee et al. 2024; Cheang, Huang et al. 2024; Prasad, Singh et al. 2024), autonomous driving (Gao, Lin et al. 2024; Hu, Wu et al. 2023; Wang, Zhao et al. 2024), and healthcare (Liu, Zhang et al. 2024). Correspondingly, the length of the videos has also increased significantly.
However, applying VLMs to long-video data introduces a significant computational challenge: The quadratic complexity of transformer-based attention mechanisms causes computation to scale rapidly with the input sequence length (Beltagy, Peters, and Cohan 2020; Vaswani et al. 2017), especially during the prefilling stage, where all input tokens must be processed before generating any output. This severely limits the scalability of VLMs in long-video understanding tasks. In practical scenarios where the entire video must be encoded before decoding begins—such as video captioning, retrieval, or reasoning—this inefficiency becomes a critical bottleneck. Therefore, optimizing the prefilling phase is essential to efficiently deploy VLMs on long video tasks.
@@ Introduction
A widely adopted technique in LLMs to speed up long context prefilling is the parallel encoding mechanism (Li et al. 2024b; Liu et al. 2024; Ma, Wang, and Tian 2025; Ratner et al. 2023; Yang, Chen, and Chen 2025; Yen, Gao, and Chen 2024; Lu et al. 2024). In these approaches, the input context is divided into blocks, and each block is encoded into Key-Value (KV) states independently. This means that tokens from different blocks do not attend to each other. Since the parallel encoding mechanism omits attention computation across blocks, it reduces the computational complexity from $O(L^{2})$ to $O(L)$. The parallel encoding mechanism also mitigates the problem of "lost in the middle" (Liu et al. 2023), and in long-context scenarios, it can even achieve accuracy surpassing that of Full-Attention (Yang, Chen, and Chen 2025).
Several state-of-the-art architectural improvements have been proposed to enhance the parallel encoding mechanism. Adaptive Parallel Encoding (APE) improves prefilling efficiency by introducing shared prefixes, an attention temperature parameter, and a scaling factor to better align the distribution of attention weight in parallel encoding with that of sequential encoding. Although these adjustments improve accuracy, they are highly sensitive to length, quantity, and content, making them difficult to tune and deploy in practice (Yang, Chen, and Chen 2025). Star Attention extends ring attention by incorporating a parallel encoding strategy, which eliminates communication overhead during the prefilling stage and reduces overall computational complexity (Liu et al. 2024). Another notable distinction of Star Attention is the use of a larger sink size, which is equal to the block size. Block-Attention & TurboRAG share the same idea: build on parallel encoding, apply position re-encoding and fint-tune to improve accuracy (Guu et al. 2020; Lu et al. 2024). However, without fine-tune, the accuracy of both suffers for the absence of a Sink Block. Nevertheless, the idea of preserving the sequential position embeddings for each block, rather than reusing a shared one, offers a valuable and insightful idea.
In addition, there are several widely discussed methods such as Native Sparse Attention (Yuan et al. 2025) from Deepseek and MoBA (Lu et al. 2025), which adopt similar algorithmic principles and achieve acceleration effects comparable to parallel encoding. However, both approaches involve modifications to the model architecture and thus require retraining from scratch. This poses a challenge for many inference platforms, which typically do not have control over users’ model preferences. The issue is particularly pronounced in domains such as autonomous driving, where a reliable models are often continuously trained and iterated over years. Consequently, we focus on finetune-free optimization methods to ensure compatibility and ease of deployment.
Given all of these, a key question naturally arises: Can existing methods be directly applied to VLMs to accelerate inference and deployment pipelines? In Table 1, we evaluate several models with existing parallel encoding methods on LongVideoBench (Zhang and Wang 2024), VideoMME (Fu et al. 2024), EgoSchema (Mangalam, Akshulakov, and Malik 2023) and MVBench (Li et al. 2024a) benchmarks. Although existing parallel encoding methods achieve some effectiveness in shorter videos, we observe varying degrees of accuracy degradation across different models on longer videos, particularly for Qwen2.5-VL. To investigate this issue, we analyze the attention weight distributions under parallel encoding and identify a misalignment between Full-Attention and parallel encoding. Existing methods developed for LLMs fail to effectively address this misalignment in VLMs. Building upon this observation. We conducted a series of studies and experiments and ultimately propose a novel attention mechanism for VLMs, termed Parallel Encoding for Vision-Language Models (PEVLM). PEVLM is a fine-tuning-free method specifically designed to accelerate the prefilling stage in long-video processing with VLMs. Our contributions involve:
  • We systematically analyze the distribution characteristics of attention weights under parallel encoding in VLMs. Our observations reveal a key reason for the misalignment between the attention weight distributions of parallel encoding and Full-Attention: the reuse of position embeddings across blocks leads to the loss of critical video information.
  • We propose PEVLM to recover the accuracy of parallel encoding by applying three targeted alignment steps: (i) Segmenting contexts into blocks by video frames rather than tokens. (ii) Using the system promots and the initial frames of the video as Sink Block for all blocks to avoid the duplication of abnormal distribution of initial tokens. (iii) Preserving the sequential position embeddings instead of resuing position embeddings across blocks. PEVLM achieves superior accuracy compared to other parallel-encoding-based methods on long video processing tasks, with up to a 9.73% improvement across different models.
  • PEVLM reduces the computational complexity of the attention mechanism from $O((T\times N)^{2})$ to $O(T\times N)$, which leading to a significant improvement in prefilling efficiency. (i) For 100k token contexts prefilling, the attention layer achieves a 7.47× speedup, while the end-to-end 2.58x speedup without compromising generation quality. This highlights the practical viability and superiority of PEVLM in cloud deployment scenarios. (ii) Under fixed latency constraints, PEVLM increases accuracy from 23.26% to 61.03%, showcasing its critical value for latency sensitive applications.
Observations
@@
(b) Parallel Encoding (Reusing Position Embedding)
(c) Parallel Encoding (Sequential Position Embedding)
@@
Figure 1: We collected the distribution of attention weights under different scenarios and computed their moving averages for better observation. The top row shows the attention weights for all tokens, while the bottom row shows those for the video tokens. (a) Full-Attention: The textual tokens and the initial video frames receive the highest attention weights, whereas the remaining video frames are assigned relatively lower attention weights. (b) Parallel Encoding (Reusing Position Embedding): The overall attention weights distribution resembles Full-Attention, with higher attention weights for the textual tokens and initial video frames. While compared to Full-Attention, the initial video frames receive even higher attention weights, and the remaining video frames exhibit more uniform and lower attention weights. (c) Parallel Encoding (Sequential Position Embedding): The overall attention weights distribution remains similar to (a) and (b). However, the attention weight distribution within the video frames more closely aligns with (a)).
To explore the application of parallel encoding in VLMs, we analyzed the potential differences between LLMs and VLMs that may affect accuracy:
(i) In VLMs, attention sinks are present not only in the initial system prompt, but also in the early video frames  (Kang et al. 2025). To achieve optimal performance, it may be necessary to include enough frames in the sink.
(ii) Existing parallel-encoding-based methods are primarily designed for Retrieval-Augmented Generation (RAG) tasks (Ma, Wang, and Tian 2025; Yang, Chen, and Chen 2025; Lu et al. 2024). In such tasks, reusing the KV cache from previously retrieved passages is crucial for enhancing inference efficiency. Since the order of passages often changes when reusing the KV cache, a common strategy is to apply parallel encoding to generate the KV cache for each passage independently, while sharing position embeddings across passages, as illustrated in Figure 3. As a result, position embeddings are typically shared across blocks. However, preserving the correct temporal order of video frames is critical for video understanding tasks. Blindly reusing position embeddings across blocks can disrupt the model’s ability to capture temporal dependencies. This disruption is especially severe for models where the temporal component of the position embeddings is explicitly aligned with absolute time (Bai et al. 2025).
Therefore, as illustrated in Figure 1, we perform an attention weight analysis for VLMs and make two key findings:
(i) As shown in Figure 1(a), initial segments of both text and video receive high attention weights, emphasizing the need to include early video frames when selecting sink tokens.
(ii) As shown in Figure 1(b), compared to Full-Attention, the initial video frames receive even higher attention weights, while the remaining frames exhibit more uniform and lower attention weights. In contrast, when sequential position embeddings are preserved, as illustrated in Figure 1(c), parallel encoding yields an attention score distribution that more closely resembles that of Full-Attention.
These observations yield two key insights for applying parallel encoding to VLMs: The shared sink block in parallel encoding should include early video frames in addition to textual prompts, as they function as important attention sinks. Maintaining the sequential position embeddings of video tokens is essential for preserving temporal consistency; reusing position embeddings across blocks can disrupt temporal alignment. These findings underscore the need for a VLM-specific parallel encoding strategy, which we introduce in the next section.
PEVLM
With the observation in the last section, we design PEVLM to improve computational efficiency. PEVLM enables a seamless shift to parallel encoding without requiring training while maintaining most of the model’s capabilities. Our approach adaptively aligns the distribution of attention weights between Full-Attention and parallel encoding, thereby boosting efficiency and performance.
Standard Attention Mechanism
@@
Attention Sink. First discovered by StreamingLLM (Lab and AI 2023), that one or more tokens at the start position of an LLM receive higher attention scores. APE (Yang, Chen, and Chen 2025) and Star-Attention (Liu et al. 2024) have demonstrated that sharing a common attention sink across parallel encoding blocks significantly enhances model accuracy. Some VLM-related studies also show that the attention sink phenomenon exists in VLMs but is more distributed among image tokens (Kang et al. 2025).
Position Embedding. To effectively process sequential input, LLMs require position embeddings, such as absolute position embeddings (Devlin et al. 2019; Vaswani et al. 2017) and relative position embeddings (Press, Smith, and Lewis 2022; Su et al. 2024). For VLMs, contextual positional information is more critical. In particular, for some VLM models like Qwen2.5-VL, their position embeddings also incorporate spatio-temporal information between video frames, making position embeddings increasingly essential for VLMs.
@@
Computational Complexity. In standard softmax attention mechanisms, the total number of operations is given by
$OP_{Attn}=2HL^{2},$ (2)
where $L$ is the context length and $H$ is the hidden size. This leads to a computational complexity of $O(L^{2})$. As $L$ increases, the cost grows quadratically, severely limiting the speed and scalability of the inference. This issue is further exacerbated in VLMs. For a video consisting of $T$ frames, where each frame is represented by $N$ tokens (typically ranging from hundreds to thousands), the total context length becomes $L=T\times N$, resulting in a complexity of $O((T\times N)^{2})$. Such scaling renders standard attention mechanisms impractical for long-video understanding tasks.
Parallel Encoding
Partitioning Strategy
As illustrated in Figure 2, our partitioning strategy consists of three core components:
@@
Sink Size & Block Size
In addition, as analyzed in the last section, the number of frames selected for the Sink Block depends primarily on the distribution of attention weights. The optimal strategy is to include all tokens with significantly higher attention scores at the beginning of the video in the Sink Block. However, as shown in Equation 11, a larger sink size results in higher latency, leading to a trade-off between accuracy and efficiency. A similar trade-off also arises in the selection of block size.
Formulations
@@
The total OP of PEVLM is
$OP_{\text{PEVLM}}=OP_{\text{Sink}}+N\times OP_{\text{ContextBlock}}+OP_{\text{Quest}},$ (6)
$OP_{\text{Sink}}=2HS^{2},$ (7)
$OP_{\text{ContextBlock}}=2HB(S+B),$ (8)
$OP_{\text{Quest}}=2HQL,$ (9)
@@
$OP_{\text{PEVLM}}=2H($ $S^{2}+Q^{2}+NB^{2}$ (10)
$+QS+NQB+NSB).$ (11)
As the block size is a static number, the computational complexity of PEVLM is $O(L)$.
Table 1: Performance ($\uparrow$) of different models and different methods on video understanding tasks evaluated at tokens from 26k to 100k.
Model Method MVBench EgoSchema VideoMME LongVideoBench
$17s\ in\ avg.$ 3 mins in avg. $<2mins$ $[2mins,1h]$ $<1min$ $[1min,1h]$
Qwen2.5-VL- Full Attn 68.47% 57.00% 74.56% 56.72% 72.85% 55.33%
7B-Instruct Block Attn 66.03% 47.40% 65.22% 50.39% 70.64% 47.85%
Max pixels: 90M APE (T=1.0) 66.03% 57.20% 50.56% 1.06% 54.02% 3.38%
Tokens: 100k Star Attn 60.19% 19.80% 3.00% 0.00% 49.03% 0.00%
PEVLM 68.44% 61.00% 74.33% 58.83% 74.52% 57.58%
LongVILA- Full Attn 61.92% 57.00% 66.33% 49.39% 65.10% 47.44%
7B-256f Block Attn 57.81% 59.00% 64.89% 51.44% 61.77% 48.05%
Frames: 256 APE (T=1.0) 60.00% 60.60% 68.56% 51.61% 63.71% 46.93%
Tokens: 66k Star Attn 63.39% 61.60% 70.11% 54.17% 63.43% 47.44%
PEVLM 63.53% 61.40% 69.44% 54.33% 65.93% 49.18%
LLaVA-Video- Full Attn 60.67% 58.20% 75.33% 59.99% 72.02% 55.53%
7B-Qwen2 (128) Block Attn 56.86% 51.60% 67.56% 55.50% 63.99% 52.36%
Frames: 128 APE (T=1.0) 58.00% 57.60% 72.22% 58.19% 67.87% 54.61%
Tokens: 26k Star Attn 59.31% 57.80% 72.89% 58.71% 68.98% 54.92%
PEVLM 60.00% 58.20% 74.33% 59.99% 72.02% 55.53%
Position Encoding
@@
As described in observations, similar to the case in LLMs, applying parallel-encoding-based methods in VLMs suffers from the misalignment of attention weight distributions when compared to Full-Attention. However, introducing additional hyperparameters to adjust attention weights for better alignment is not ideal for efficient deployment. Therefore, inspired by Block Attention (Ma, Wang, and Tian 2025) and TurboRAG (Lu et al. 2024), we retain sequential position embeddings as illustrated in Figure 3.
Due to retaining the sequential position embeddings, PEVLM does not extend the model’s inherent capacity to handle longer video contexts, unlike some other parallel encoding-based approaches. For instance, LongVILA-7B-256f is generally trained to process up to 256 frames, and feeding longer videos may lead to accuracy degradation. Although PEVLM accelerates inference, it still processes 256 frames, and longer inputs may still result in degraded performance. However, this is acceptable, as our primary objective is to improve prefilling efficiency.
Experiments